# Euler-type approximation for the invariant measure: An abstract framework

### v1

@@ B Moreover $\theta$ is $(2,b_{*},1)$-self-coupled for some $h_{0}>0$ and more precisely
By (B.17), ${\mathbb{E}}\left|x_{s,r}^{2}-X^{2}\right|^{2}\leq C(1+{\mathbb{E}}\left|X^{2}\right|)(r-s).$ We also recall that by the very definition of $\tau$, the inequality $2xy\leq\delta x^{2}+\delta^{-1}y^{2}$ and (B.17), we have
$\int_{0}^{1}\left|\tau_{s,r}^{1}(w)-\tau_{s,r}^{2}(w)\right|^{2}dw$ $=$ $W_{2}^{2}(\Theta^{m}_{s,r}(\mu^{1}),\mu^{2})\leq{\mathbb{E}}\left(|x_{s,r}^{1}(X^{1})-X^{2}|^{2}\right)$
$\leq$ $(1+\delta){\mathbb{E}}(|X^{1}-X^{2}|^{2})+(1+{\delta}^{-1}){\mathbb{E}}\left(|x_{s,r}^{1}-X^{1}|^{2}\right)$
$\leq$ $(1+\delta){\mathbb{E}}(|X^{1}-X^{2}|^{2})+C(1+{\delta}^{-1})(1+{\mathbb{E}}(|X^{1}|^{2}))(t-s).$
@@
${\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\left|\Delta c(r,w,z)\right|^{2}dw\nu(dz)dr$ (B.23)
$\leq$ ${\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\overline{c}^{2}(z)(\left|\tau_{s,r}^{1}(w)-\tau_{s,r}^{2}(w)\right|^{2}+\left|x_{s,r}^{1}-X^{2}\right|^{2})dw\nu(dz)dr$
$\leq$ $2Q(1+\delta){\mathbb{E}}(|X^{1}-X^{2}|^{2})+C(1+{\delta}^{-1})(1+{\mathbb{E}}(|X^{1}|^{2}))(t-s).$
Then, the above estimates in (B.22) and (B.23) allow us to take expectations in the Itô’s formula for the function $x\rightarrow\left|x\right|^{2}$. We also use the identity $\left|y+\Delta c\right|^{2}-\left|y\right|^{2}-2\left\langle y,\Delta c\right\rangle=\left|\Delta c\right|^{2},$ so that we obtain
${\mathbb{E}}(|y_{s,t}|^{2})$ $=$ ${\mathbb{E}}(|X^{1}-X^{2}|^{2})+2{\mathbb{E}}\int_{s}^{t}\left\langle y_{s,r},\Delta b(r)\right\rangle dr+{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\left|\Delta c(r,w,z)\right|^{2}dw\nu(dz)dr.$
$\leq$ ${\mathbb{E}}(|X^{1}-X^{2}|^{2})(1+2(Q+\delta(2+Q)-\overline{b})(t-s))+C(1+{\mathbb{E}}(|X^{1}|^{2})+{\mathbb{E}}(|X^{2}|^{2}))(t-s)^{3/2}$
so (B.20) is proved by choosing $\delta$ small enough. In particular, since ${\mathbb{E}}\left|X^{1}-X^{2}\right|^{2}=W_{2}^{2}(\mu^{1},\mu^{2}),$ this gives
@@
$\leq$ ${\mathbb{E}}\left(|x_{s,t}^{1}-x_{s,t}^{2}|^{2}\right)={\mathbb{E}}(|y_{s,t}|^{2})$
$\leq$ $W_{2}^{2}(\mu^{1},\mu^{2})(1+2(Q+\delta(2+Q)-\overline{b})(t-s))+C(1+\left\|\mu^{1}\right\|_{2}^{2}+\left\|\mu^{2}\right\|_{2}^{2})(t-s)^{3/2}$
Now, using (5.39) and taking the limit as $m\rightarrow\infty$ we get

### v2

@@ B Moreover $\theta$ is $(2,b_{*},1)$-self-coupled for some $h_{0}>0$ and more precisely
By (C.20), ${\mathbb{E}}[|x_{s,r}^{2}-X^{2}|^{2}]\leq C(1+{\mathbb{E}}[|X^{2}|])(r-s).$ We also recall that by the very definition of $\tau$, the inequality $2xy\leq\delta x^{2}+\delta^{-1}y^{2}$ and (C.20), we have
$\!\!\int_{0}^{1}\!\!\left|\tau_{s,r}^{1}(w)-\tau_{s,r}^{2}(w)\right|^{2}\!dw\!$ $=W_{2}^{2}(\Theta^{m}_{s,r}(\mu^{1}),\mu^{2})\leq{\mathbb{E}}\left[|x_{s,r}^{1}(X^{1})-X^{2}|^{2}\right]$
$\leq(1+\delta){\mathbb{E}}[|X^{1}-X^{2}|^{2}]+(1+{\delta}^{-1}){\mathbb{E}}\left[|x_{s,r}^{1}-X^{1}|^{2}\right]$
$\leq(1+\delta){\mathbb{E}}[|X^{1}-X^{2}|^{2}]+C(1+{\delta}^{-1})(1+{\mathbb{E}}[|X^{1}|^{2}])(t-s).$
@@
${\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\left|\Delta c(r,w,z)\right|^{2}dw\Lambda(dz)dr$ (C.26)
$\leq$ ${\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\overline{c}^{2}(z)(\left|\tau_{s,r}^{1}(w)-\tau_{s,r}^{2}(w)\right|^{2}+\left|x_{s,r}^{1}-X^{2}\right|^{2})dw\Lambda(dz)dr$
$\leq$ $2Q(1+\delta){\mathbb{E}}[|X^{1}-X^{2}|^{2}]+C(1+{\delta}^{-1})(1+{\mathbb{E}}[|X^{1}|^{2}])(t-s).$
Then, the above estimates in (C.25) and (C.26) allow us to take expectations in the Itô’s formula for the function $x\rightarrow\left|x\right|^{2}$. We also use the identity $\left|y+\Delta c\right|^{2}-\left|y\right|^{2}-2\left\langle y,\Delta c\right\rangle=\left|\Delta c\right|^{2},$ so that we obtain
${\mathbb{E}}[|y_{s,t}|^{2}]=$ ${\mathbb{E}}[|X^{1}-X^{2}|^{2}]+2{\mathbb{E}}\int_{s}^{t}\left\langle y_{s,r},\Delta b(r)\right\rangle dr$
$+{\mathbb{E}}\int_{s}^{t}\int_{E}\int_{0}^{1}\left|\Delta c(r,w,z)\right|^{2}dw\Lambda(dz)dr.$
$\leq$ ${\mathbb{E}}[|X^{1}-X^{2}|^{2}](1+2(Q+\delta(2+Q)-\overline{b})(t-s))$
$+C(1+{\mathbb{E}}[|X^{1}|^{2}]+{\mathbb{E}}[|X^{2}|^{2}])(t-s)^{3/2}$
so (C.23) is proved by choosing $\delta$ small enough. In particular, since ${\mathbb{E}}[|X^{1}-X^{2}|^{2}]=W_{2}^{2}(\mu^{1},\mu^{2}),$ this gives
@@
$\leq{\mathbb{E}}\left[|x_{s,t}^{1}-x_{s,t}^{2}|^{2}\right]={\mathbb{E}}[|y_{s,t}|^{2}]$
$\leq W_{2}^{2}(\mu^{1},\mu^{2})(1+2(Q+\delta(2+Q)-\overline{b})(t-s))+C(1+\left\|\mu^{1}\right\|_{2}^{2}+\left\|\mu^{2}\right\|_{2}^{2})(t-s)^{3/2}.$
Now, using (C.16) and taking the limit as $m\rightarrow\infty$ we get
@@
Acknowledgment: A.A. acknowledges the support of the ”chaire Risques Financiers”, Fondation du Risque.